write "t" as a function of "Q"


Q(t) = Qo ( 1-e^-t/a)

Answer by stanbon(75887)

write "t" as a function of "Q"
---------------------
Q(t) = Qo ( 1-e^-t/a)
----
(1 - e^(-t/a)) = Q(t)/Qo
=====
e^(-t/a) = 1-[Q(t)/Qo]
------
-t/a = ln[1-Q(t)/Qo ]
------
t = -a*ln[1-Q(t)/Qo]
-------------
Cheers,
Stan H.
----------------